OpenStudy (anonymous):

Factor completely: 5ab + 3ay + 5b + 3y

OpenStudy (anonymous):

5ab+3ay+5b+3y =a(5b+3y)+1(5b+3y) =(5b+3y)(a+1)

OpenStudy (hoblos):

= a(5b+3y) + 5b+3y = (5b+3y)(a+1)